OpenStudy (anonymous):

I'm blooped. Fruit juice makes up 2/3 of the liquid ingredients in Alonzo's punch recipe. Estimate how many gallons are needed to make 8 1/2 gallons of punch. Thanks

OpenStudy (triciaal):

i would use ratio fruit juice:liquid = 2/3 let amount of fruit juice = x ; amount of liquid made 8 1/2 = 17/2 x/(17/2) = 2/3 solve for x

OpenStudy (triciaal):

17/3 = 5 2/3 gal